OpenStudy (anonymous):

last valentine's day 100 red roses cost r dollars. this valentine's day,80 of the same kind of rose cost .2r dollars. if there were no discounts based on the size of the purchase, what was the percentage decrease in the cost of a red rose?

OpenStudy (anonymous):

100 roses cost r then each rose costs 100/r 80 roses cost .2r then each costs 80/.2r To get the percentage decrease we use this formula: (New - Old)/Old or [80/.2r - 100/r]/(100/r)

OpenStudy (anonymous):

@seraphic_topaz i got 300%.......i still need help

OpenStudy (anonymous):

Let's simplify the numeral: 80/.2r - 100/r = (80 - 100*.2)/.2r = (80 - 20)/.2r = 60/.2r [60/.2r]/[100/r] = [60/.2r]*[r/100] =60/[.2*100] = 60/20 = 3. Then you must be correct.

OpenStudy (anonymous):

so when u convert 3 to percentage it is 300%

OpenStudy (anonymous):

OOPS! I did it again. Why did I do 100/r or 80/.2r?? It should have been r/100 and .2r/80. And then you can apply the New-Old formula. My bad.

OpenStudy (anonymous):

100 roses cost r then each rose costs r/100; 80 roses cost .2r then each rose costs .2r/80. [.2r/80 - r/100]/[r/100] should give you the answer.

OpenStudy (anonymous):

uhhh 75%

OpenStudy (anonymous):

.75 < 1 so this answer is more reasonable than the last one because we're looking for percentage decrease.